# Adaptive Agents in Spatial Double-Auction Markets: Modeling the Emergence of Industrial Symbiosis

### v1

@@ 3.4. Symbiosis Index
$\text{SI}=\frac{q_{\text{bought}}}{\min({q_{\text{toSell}},q_{\text{needed}}})},$
where $q_{\text{bought}}$ denotes the quantity actually purchased, $q_{\text{toSell}}$ the quantity offered by the seller, and $q_{\text{needed}}$ the quantity required by the buyer. The denominator normalizes the indicator so that it cannot exceed 1, even when the sold quantity is greater than the buyer’s need. This ensures that the symbiosis index reflects the proportion of demand satisfied rather than rewarding oversupply. In contrast to existing technical indicators hardy2002industrial; tiejun2010two; zhao2008simulation; gao2013study, which often incorporate numerous parameters (e.g., material compatibility, lifecycle impacts, or multi-criteria environmental scores), our simplified formulation deliberately focuses only on exchanged quantities. This avoids unnecessary complexity while remaining sufficient for capturing efficiency in the present modeling framework.
@@ 4.1. Simulation Framework
The framework tracks a variety of systemic metrics over time, including:
  • Local prices and their evolution;
  • Proportion of demand satisfied locally (symbiosis index);
  • Network connectivity and trading patterns;
  • Distance travelled by the sold byproducts.
These indicators form the backbone of our analysis, as they connect local agent decisions to measurable systemic dynamics. This flexibility enables us to investigate a wide range of experimental settings, from dense and highly interconnected territories to sparse environments where exchanges are scarce. Visualization tools provide spatial and temporal representations of market dynamics, price formation, and circularity indicators. The code implementation is available on GitLab [^1] https://github.com/AAMAS746/aamas2026-746.
4.2. Experimental Conditions.
For similar reasons behind our decision to focus on a single product, we refrain from using a real-world territory and instead construct virtual environments. This approach provides precise control over key parameters such as firm density, territorial extent, and spatial clustering, thereby enabling a systematic exploration of their influence on circularity indicators and overall market performance. One of the key factors shaping the potential for cooperation is the spatial dispersion of firms. We therefore define the cluster spread parameter ($cs$) as the standard deviation of companies positions around cluster centers, normalized by the environment width: small values lead to compact, well separated clusters, while setting it to $1$ produces a configuration comparable to a fully uniform distribution across the environment, effectively removing any clustering effect. Figure 1 shows the difference between a low (1(a)) and high (1(b)) spread value. In our experiments, businesses were allocated into four spatial clusters. By default, the cluster spread parameter was set to $1$, indicating a broad dispersion and minimal clustering.
@@
The results we obtain came from simulations conducted with a population of 40 firms. The market price was fixed at $100$, and the transportation cost was set to $0.1$ per kilometer. Sellers adjusted their strategies through Boltzmann exploration, with a temperature decay parameter set to $0.996$ and the action space discretized into $30$ bins. The relatively small scale of the simulated market was chosen to balance computational tractability with the ability to capture complex multi-agent interactions. This size of population allows us to conduct thousands of simulation runs for systematic sensitivity analyses and counterfactual experiments, which would be computationally prohibitive at larger scales. Furthermore, each simulation runs for a large number of timesteps (1000 steps) to ensure that seller strategies have sufficient time to adapt and converge, and that emergent patterns of price formation and allocation stabilize. By prioritizing repeated experimentation and long-horizon dynamics over absolute market size, we focus on understanding the fundamental mechanisms driving spatial effects, adaptive bidding behavior, and the emergence of circularity. Our implementation is optimized for performance, allowing on a standard laptop processor (Intel(R) Core(TM) Ultra 9 185H processor with 21 cores) to achieve approximately 100 simulations of 1000 timesteps per minute.
@@ 5.1. Convergence Toward Equilibrium
Figure 2. Evolution of mean transaction prices under different scarcity and disposal cost conditions.
@@
Let $\phi_{j}^{(t)}\in\Phi_{j}$ denote the action used by seller $s$ at timestep $t$, and $r_{j}^{(t)}(\phi_{j}^{(t)})$ denote the reward received at that step. The optimal reward at time $t$ is
@@
where $r_{j}^{(t)}(\phi)$ is the reward seller $s$ would have obtained at step $t$ if it had chosen action $\phi$ while all other sellers’ actions remained unchanged. The regret of seller $s$ at time $t$ is then
@@
Empirically, we observe that per-step total regret, the aggregate deviation from a joint optimum, declines over time and converges toward zero in both scenarios (Figure 3). Since per-step regret measures the immediate payoff loss relative to the best fixed action in hindsight at that step, a vanishing regret implies that sellers’ strategies are nearly optimal given the current actions of others. Consequently, the joint strategy profile of all sellers appears to converge toward an approximate Nash equilibrium, providing justification for the stability of the emergent market dynamics under our MARL learning scheme.
@@ 5.2. Price and Symbiosis Dynamics
In this section, we examine how firm density, resource scarcity, and disposal costs shape the dynamics of transaction prices and circularity. We focus on these output metrics because they directly quantify the level of local circularity achieved in the market and the average exchange price, making them relevant proxies for assessing the effectiveness of different policy levers. By systematically varying firm density, resource scarcity, and disposal costs, we analyze how structural and economic factors influence sellers’ adaptive strategies, shaping both the emergence of local symbiosis and overall market outcomes. Figures 4 and 5 illustrate the evolution of equilibrium prices and local symbiosis levels as a function of the disposal cost, the scarcity, and the density. Each point is the mean of 10 simulation runs, with shaded areas indicating one standard deviation to visualize variability across runs.
@@
Symbiosis levels (Figure 5) exhibit a complementary pattern: the share of demand met through local exchanges increases with $c_{d}$, as the disposal penalty strengthens the incentive to form local contracts. Here, density plays a critical role: in dense environments, local exchanges dominate rapidly and symbiosis saturates close to one; in sparse environments, by contrast, symbiosis remains low even under high $c_{d}$, reflecting structural limitations in firms’ connectivity. We observe that this time scarcity has only a marginal effect on the evolution of symbiosis: higher scarcity slightly enhances local exchanges, but the dominant factors remain disposal cost and firm density.
@@ 5.3. Sensitivity Analysis
The previous analysis showed that market outcomes are shaped by interacting mechanisms, but it did not allow us to fully disentangle their relative influence. To address this limitation and account for the influence of additional variables, we conducted a variance-based global sensitivity analysis saltelli2010variance. We evaluate the influence of the simulation parameters on the symbiosis index and converged mean price measured at the end of each simulation run, after 1000 timesteps. By analysing which parameters exert the greatest influence on the price and symbiosis index, we can identify the factors that policymakers should prioritize when designing interventions to foster circularity while keeping prices at a low level. We simulate for varying values of the five following variables:
@@
  • Scarcity level ($s\in[0.25,2]$), controlling the relative abundancy of resources in the market;
@@
To efficiently explore the parameter space, we employed surrogate models, which act as fast emulators of the full simulation while capturing the key input-output relationships saves2024smt. Specifically, we implemented a sparse Polynomial Chaos Expansion (PCE), enabling the analytical derivation of Sobol’ indices from its coefficients sudret2008global. The surrogate models demonstrated strong predictive performance, achieving an $R^{2}$ of 95% for the symbiosis and 92% for the price on a 20% test sample. Parameter sampling was performed using Latin hypercube sampling to train surrogate models, with $s$ sampled on a $\log_{2}$ scale and density{density} on a $\log_{10}$ scale. For each parameter set, two independent simulations were executed to account for stochastic variability in agent decisions, resulting in a total of 20,000 independent simulations.
Sobol’ sensitivity analysis decomposes the variance of the model output into orthogonal contributions attributable to individual input parameters and their interactions. We computed the full set of Sobol’ indices: first-order indices quantify the direct effect of each parameter on output variance, while higher-order indices capture interaction effects amongst parameters chastaing. Taken together, Sobol’ indices provide a rigorous and comprehensive assessment of the relative importance of independent inputs and their interactions, with their sum accounting for $100\%$ of the variance.
@@
Even if higher density is highly influential for both our symbiosis and price objectives, in practice, density is constrained by the considered territory and is not directly actionable. Therefore, in the subsequent analysis, we focus on the 4 other input variables under different fixed density scenarios, since their interaction effects with density are all significant even at a third order. To better capture the non-linear effects, we replace the sparse PCE by a Multi-Layer Perceptron neural network surrogate model ($R^{2}$ of 97% for the symbiosis and 95% for the price), and use it to predict both the Partial Dependence Plots (PDP) and Individual Conditional Expectations (ICE) robani2025smt as illustrated in Figure 6. In these plots, each colored line corresponds to an ICE, showing how the outcome of a single simulation responds when varying one parameter while holding others fixed; the black line is the PDP, representing the average effect across all simulations, and the color of each ICE line encodes the underlying firm density.
@@
(a) PDP and ICE for the price index
(b) PDP and ICE for the symbiosis index
@@
In summary, firm density is a primary determinant of both symbiosis and price, though it is mostly exogenous and difficult to influence directly through policy. Amongst the controllable parameters, transport cost ($c_{t}$) and disposal penalty ($c_{d}$) emerge as effective levers. The sensitivity results indicate that reducing per-kilometer transport costs, through measures such as shared logistics or targeted subsidies, can substantially enhance exchange opportunities in low-density areas. Similarly, higher disposal penalties consistently incentivize firms to seek symbiotic exchanges instead of wasteful disposal.
@@ 6. Discussion
Our simulations highlight several insights into the dynamics of decentralized byproduct markets. First, the emergent pricing behavior shows that the RL process enables sellers to adapt efficiently to local conditions. The counterfactual regret analysis further demonstrates that the learned strategies converge toward a near Nash equilibrium, suggesting that decentralized adaptive learning can yield stable and rational outcomes without requiring a centralized market clearing mechanism. This finding underscores the potential of bottom-up coordination mechanisms in fostering circular exchanges.
The symbiosis index provides a tractable measure of local circularity and illustrates how territorial factors, scarcity, kilometer and disposal costs shape resource exchanges. Our results show that transportation costs and disposal penalties represent actionable levers, especially where density is not naturally high. Moreover, interactions among variables, including up to third-order effects, highlight that circularity outcomes result from the combined influence of spatial configuration and economic incentives rather than any single factor. These findings emphasize the importance of integrated, multivariate approaches when designing circular economy interventions.
Beyond methodological insights, the model offers a structured framework for evaluating policy instruments. The parameters of the simulation naturally correspond to concrete levers available to regulators. Increasing the disposal cost $c_{d}$ can be interpreted as a landfill tax, designed to encourage firms to recover or valorize byproducts locally instead of discarding them. Adjusting the external market price $p_{m}$ through tariffs or subsidies can favor local sourcing over reliance on virgin raw materials. Similarly, incorporating carbon related transportation costs km_costkm\_cost provides a means to internalize the environmental footprint of long distance exchanges. Finally, promoting spatial proximity through industrial zoning or cluster based planning can strengthen firm density, thereby enhancing the feasibility of symbiotic interactions.
While these policy-relevant insights highlight the practical applicability of our model, it is important to recognize several limitations that temper the generalizability of the results. The present model focuses on a single byproduct, which allows us to isolate core mechanisms of decentralized exchange but leaves aside interdependencies across multiple byproduct markets. Real-world industrial symbiosis often involves multi-product complementarities, cascading exchanges, and competition between alternative uses of resources. Likewise, our experiments relied on stylized virtual geographies rather than empirically calibrated territories. While this approach ensured control and replicability, it abstracts away from infrastructural constraints such as road networks or administrative boundaries that can strongly affect exchange feasibility. Another simplification concerns the temporal dimension: in our simulations, companies participate in auctions independently of their production cycles, whereas in practice production schedules and storage constraints strongly influence availability and willingness to trade. Furthermore, our agents trade without any memory of past interactions, while in reality firms build relationships over time, developing trust, reliability, and long term contracting practices. Nonetheless, by abstracting from these complexities, the present work provides an indispensable first step toward systematically modeling spatially constrained markets for industrial symbiosis, upon which richer and more realistic frameworks can be built.
These limitations also point to promising directions for future work. Extending the model to multi-product environments would allow capturing substitution effects and synergies between resource flows. Introducing temporal dynamics and agent memory would bring the model closer to real-world conditions, enabling the study of production rhythms, storage capacities, and relational contracts. Grounding the simulations in empirical case studies of specific territories would make it possible to validate predictions and directly inform policy design. Ultimately, developing a panel of decision parameters that systematically map simulation outcomes to actionable policies could provide decision makers with a robust tool to design interventions that align individual incentives with collective circular economy objectives.
@@ 7. Conclusion
We introduced a decentralized multi-agent model for local byproduct exchanges, capturing how sellers adapt prices through RL, based on spatial and economic constraints. Simulations exhibit that adaptive decentralized strategies can produce stable, efficient, and circular market outcomes, as evidenced by low counterfactual regret and high symbiosis indices. Our results highlight two complementary contributions.
First, from a policy design perspective, the simulator is the first step toward a decision support tool. By linking agent-level parameters to macro-level outcomes such as the symbiosis index, the model allows policymakers to identify which levers most strongly affect circularity. This provides a controlled environment where different regulatory scenarios can be tested before real-world implementation. The emergent insights from our simulations help policymakers prioritize interventions that effectively promote circular practices across heterogeneous firms and contexts.
Second, from a MAS perspective, the work provides a new methodological understanding of decentralized exchange mechanisms under spatial and economic constraints. We show how RL agents, interacting in a double auction market with heterogeneous resources and transport costs, can converge to stable trading patterns. Beyond the application domain, our contribution lies in revealing how market based coordination and learning dynamics shape emergent equilibria.
@@
This work is part of the MUTTEC project, funded by the France 2030 program – TIRIS (Interdisciplinarité et impact de la science pour la société), under contract number 23-AAP-TIRIS-01-062. This work is part of the MIMICO research project funded by the Agence Nationale de la Recherche (ANR) no ANR-24-CE23-0380.

### v2

@@ 3.4. Symbiosis Index
$\text{SI}=\frac{q_{\text{bought}}}{\min({q_{\text{toSell}},q_{\text{needed}}})},$ (7)
where $q_{\text{bought}}$ denotes the total quantity actually purchased, $q_{\text{toSell}}$ the total quantity offered by the sellers, and $q_{\text{needed}}$ the total quantity required by the buyers. The denominator normalizes the indicator so that it cannot exceed 1, even when the sold quantity is greater than the buyer’s need. This ensures that the symbiosis index reflects the proportion of demand satisfied rather than rewarding oversupply. In contrast to existing technical indicators hardy2002industrial; tiejun2010two; zhao2008simulation; gao2013study, which often incorporate numerous parameters (e.g., material compatibility, life cycle impacts, or multi-criteria environmental scores), our simplified formulation deliberately focuses only on exchanged quantities. This avoids unnecessary complexity while remaining sufficient for capturing efficiency in the present modeling framework.
@@ 4.1. Simulation Framework
The framework tracks a variety of systemic metrics over time, including the local prices and their evolution, the proportion of demand satisfied locally (symbiosis index) and the trading patterns. These indicators form the backbone of our analysis, as they connect local agent decisions to measurable systemic dynamics. This flexibility enables us to investigate a wide range of experimental settings, from dense and highly interconnected territories to sparse environments where exchanges are scarce. Visualization tools provide spatial and temporal representations of market dynamics, price formation, and circularity indicators. The code implementation is open source [^1] https://github.com/matthieu-mastio/MASTIO.
4.2. Experimental Conditions
For similar reasons behind our decision to focus on a single product, we decided to focalize on virtual environments. This approach provides precise control over key parameters such as firm density, territorial extent, and spatial clustering, thereby enabling a systematic exploration of their influence on circularity indicators and overall market performance. One of the key factors shaping the potential for cooperation is the spatial dispersion of firms. We therefore define the cluster spread parameter ($cs$) as the standard deviation of companies’ positions around cluster centers, normalized by the environment width: small values lead to compact, well-separated clusters, while setting it to $1$ produces a configuration comparable to a fully uniform distribution across the environment, effectively removing any clustering effect. Figure 1(b) shows the difference between a low (1(a)) and high (1(b)) spread value. In our experiments, businesses were allocated into four spatial clusters. By default, the cluster spread parameter was set to $1$, indicating a broad dispersion and minimal clustering.
@@ 4.2. Experimental Conditions
[…]
Comparison of spatial agent distributions for different cluster spread values: 2 cases are displayed, the first one with cs = 0.05, showing 4 dense clusters; the second one with cs = 0.5, displaying an almost uniform distribution of agents in space.
The results we obtained came from simulations conducted with a population of 40 firms, representing the typical scale of a standard EIP. The market price was fixed at $100$, and the transportation cost was set to $0.1$ per kilometer. This transport cost is a normalized parameter relative to the market price, calibrated to ensure that spatial distance acts as a meaningful economic friction without completely blocking trade, a choice further supported by the sensitivity analysis in Section 5.3. Sellers adjusted their strategies through Boltzmann exploration, with a temperature decay parameter set to $0.996$ and the action space discretized into $30$ bins. The relatively small scale of the simulated market was chosen to balance computational tractability with the ability to capture complex multi-agent interactions. This size of population allows us to conduct thousands of simulation runs for systematic sensitivity analyses and counterfactual experiments, which would be computationally prohibitive at larger scales. Furthermore, each simulation runs for a large number of timesteps (1000 steps) to ensure that seller strategies have sufficient time to adapt and converge, and that emergent patterns of price formation and allocation stabilize. By prioritizing repeated experimentation and long-horizon dynamics over absolute market size, we focus on understanding the fundamental mechanisms driving spatial effects, adaptive bidding behavior, and the emergence of circularity. Our implementation is optimized for performance, allowing on a standard laptop processor (Intel(R) Core(TM) Ultra 9 185H processor with 21 cores) to achieve approximately 100 simulations of 1000 timesteps per minute.
@@ 5.1. Convergence Toward Equilibrium
Figure 2. Evolution of mean transaction prices under different scarcity and disposal cost conditions, where $s$ denotes the scarcity ratio and $c_{d}$ the disposal cost.
@@
Let $\phi_{j}^{(t)}\in\Phi_{j}$ denote the action used by seller $j$ at timestep $t$, and $r_{j}^{(t)}(\phi_{j}^{(t)})$ denote the reward received at that step. The optimal reward at time $t$ is
@@
where $r_{j}^{(t)}(\phi)$ is the reward that the seller $j$ would have obtained at step $t$ if it had chosen action $\phi$ while all other sellers’ actions remained unchanged. The regret of seller $j$ at time $t$ is then
@@
Empirically, we observe that the per-step total regret, the aggregate deviation from a joint optimum, declines over time and converges toward zero in both scenarios (Figure 3). Since per-step regret measures the immediate payoff loss relative to the best fixed action in hindsight at that step, a vanishing regret implies that sellers’ strategies are nearly optimal given the current actions of others. Consequently, the joint strategy profile of all sellers appears to converge toward an approximate Nash equilibrium, providing justification for the stability of the emergent market dynamics under our MARL learning scheme.
@@ 5.2. Price and Symbiosis Dynamics
In this section, we examine how firm density, resource scarcity, and disposal costs shape the dynamics of transaction prices and circularity. We focus on these output metrics because they directly quantify the level of local circularity achieved in the market and the average exchange price, making them relevant proxies for assessing the effectiveness of different policy levers. By systematically varying firm density, resource scarcity, and disposal costs, we analyze how structural and economic factors influence sellers’ adaptive strategies, shaping both the emergence of local symbiosis and overall market outcomes. Figures 4 and 5 illustrate the evolution of equilibrium prices and local symbiosis levels as a function of disposal cost, scarcity, and density. Each point is the mean of 10 simulation runs, with shaded areas indicating one standard deviation to visualize variability across runs.
@@
Symbiosis levels (Figure 5) exhibit a complementary pattern: the share of demand met through local exchanges increases with $c_{d}$, as the disposal penalty strengthens the incentive to form local contracts. Here, density plays a critical role: in dense environments, local exchanges dominate rapidly, and symbiosis saturates close to one; in sparse environments, by contrast, symbiosis remains low even under high $c_{d}$, reflecting structural limitations in firms’ connectivity. We observe that this time scarcity has only a marginal effect on the evolution of symbiosis: higher scarcity slightly enhances local exchanges, but the dominant factors remain disposal cost and firm density.
@@ 5.3. Sensitivity Analysis
The previous analysis showed that market outcomes are shaped by interacting mechanisms, but it did not allow us to fully disentangle their relative influence. To address this limitation and account for the influence of additional variables, we conducted a variance-based global sensitivity analysis saltelli2010variance.
We evaluate the influence of the simulation parameters on the symbiosis index and exchanged prices measured at the end of each simulation run after 1000 timesteps. By analyzing which parameters exert the greatest influence on the price and symbiosis index, we can identify the factors that policymakers should prioritize when designing interventions to foster circularity while keeping prices at a low level. We simulate varying values for the following five variables:
@@
  • Scarcity level ($s\in[0.25,2]$), controlling the relative abundance of resources in the market;
@@
To efficiently explore the parameter space, we employed surrogate models, which act as fast emulators of the full simulation while capturing the key input-output relationships saves2024smt. Specifically, we implemented a sparse Polynomial Chaos Expansion (PCE), enabling the analytical derivation of Sobol’ indices from its coefficients sudret2008global. The surrogate models demonstrated strong predictive performance, achieving an $R^{2}$ of 95% for the symbiosis and 92% for the price on a 20% test sample. Parameter sampling was performed using Latin hypercube sampling to train surrogate models, with $s$ sampled on a $\log_{2}$ scale and density{density} on a $\log_{10}$ scale. For each parameter set, two independent simulations were executed to account for stochastic variability in agent decisions, resulting in a total of 20,000 independent simulations. Sobol’ sensitivity analysis decomposes the variance of the model output into orthogonal contributions attributable to individual input parameters and their interactions. We computed the full set of Sobol’ indices: first-order indices quantify the direct effect of each parameter on output variance, while higher-order indices capture interaction effects among parameters chastaing. Taken together, Sobol’ indices provide a rigorous and comprehensive assessment of the relative importance of independent inputs and their interactions, with their sum accounting for $100\%$ of the variance.
@@
Although firm density emerges as the most influential factor for both symbiosis and price, its overwhelming effect tends to dominate the variance and obscure the individual contributions of the other parameters. To better isolate and interpret the impact of the potentially actionable economic and spatial levers, we therefore fix density at different levels and focus the subsequent analysis on the remaining four input variables. Therefore, in the subsequent analysis, we focus on the 4 other input variables under different fixed density scenarios, since their interaction effects with density are all significant even at a third order. To better capture the non-linear effects, we replace the sparse PCE by a Multi-Layer Perceptron neural network surrogate model ($R^{2}$ of 97% for the symbiosis and 95% for the price), and use it to predict both the Partial Dependence Plots (PDP) and Individual Conditional Expectations (ICE) robani2025smt as illustrated in Figure 5.3. In these plots, each colored line corresponds to an ICE, showing how the outcome of a single simulation responds when varying one parameter while holding others fixed; the black line is the PDP, representing the average effect across all simulations, and the color of each ICE line encodes the underlying firm density.
@@
(a) PDP and ICE illustrating the factors influencing the price index
(b) PDP and ICE illustrating the factors influencing the symbiosis index
@@
[…]
@@ 6. Discussion
Our results demonstrate that the proposed mechanism generates stable, economically grounded behaviors. Emergent pricing patterns reveal that sellers adapt efficiently to local conditions, while counterfactual regret analysis confirms convergence toward a near-Nash equilibrium rather than stochastic oscillation. Furthermore, global sensitivity analysis establishes clear causal links between model parameters and market outcomes. Significant high-order interactions suggest that circularity arises from the synergy of spatial configuration and economic incentives, highlighting the necessity for integrated, multivariate policy frameworks. Ultimately, these findings indicate that decentralized adaptive learning produces rational outcomes without centralized clearing, validating the mechanism’s robustness. Beyond methodology, the model offers a framework for evaluating policy instruments by mapping simulation parameters to regulatory levers. For instance, increasing disposal costs mimics landfill taxes, while the cluster spread parameter can be interpreted as a proxy for industrial zoning or policies that promote firm density. This mapping enables the simulator to function as a virtual laboratory, allowing stakeholders to perform “what-if” analyzes and compare interventions before real-world deployment. While these insights highlight the applicability of our model, we must recognize several limitations that temper the generalizability of the results. We focused on a single byproduct to isolate the core mechanisms of decentralized exchange, abstracting away product heterogeneity and multi-market interdependence. Likewise, to ensure control and replicability, our experiments relied on stylized virtual geographies rather than empirically calibrated territories. We also abstract from temporal dynamics (production cycles, storage) and agent memory (trust, long-term contracts). Nonetheless, this work provides an indispensable first step toward systematically modeling spatially constrained markets for industrial symbiosis, upon which richer and more realistic frameworks can be built. Future work will address these limitations by introducing multi-product environments, temporal dynamics, and agent memory to capture substitution effects and relational contracts. Grounding simulations in empirical case studies will further validate predictions and refine the policy decision tool. Finally, we also plan to include rigorous runtime benchmarking to formally showcase the simulator’s scalability.
@@ 7. Conclusion
We introduced a decentralized multi-agent model for local byproduct exchanges, capturing how sellers adapt prices through RL, based on spatial and economic constraints. Simulations exhibit that adaptive decentralized strategies can produce stable circular outcomes. Our results highlight two complementary contributions. First, from a policy design perspective, the simulator is the initial step toward a decision support tool. By linking agent-level parameters to macro-level outcomes such as the symbiosis index, the model allows policymakers to identify which levers most strongly affect circularity. This provides a controlled environment where different regulatory scenarios can be tested before real-world implementation. The emergent insights from our simulations help policymakers prioritize interventions that effectively promote circular practices across heterogeneous firms and contexts. Second, from a MAS perspective, the work provides a new methodological understanding of decentralized exchange mechanisms under spatial and economic constraints. We show how RL agents, interacting in a double auction market with heterogeneous resources and transport costs, can converge to stable trading patterns. Beyond the application domain, our contribution lies in revealing how market based coordination and learning dynamics shape emergent equilibrium.
@@
This work was supported by the MUTTEC project (France 2030 – TIRIS, contract 23-AAP-TIRIS-01-062) and by the MIMICO project (ANR-24-CE23-0380).